# Wally: An Efficient Private Search Engine

### v3

@@ 4.2 SHE details in keyword search
At a high level, for each bucket independently, the server uses the cuckoo hashing to map the $N^{\prime}$ entries into a table of size $O(N^{\prime})$ using two random hash functions. The guarantee is that if a keyword is present in the database, it must be at bucket index $h_{1}(\textit{keyword})$ or $h_{2}(\textit{keyword})$ in the table. The client then uses index PIR to privately fetch entries at these indices and locally find the index containing the entry and associated data. Wally uses MulPIR [4] as the underlying index PIR, an efficient single-server PIR scheme based on SHE. We emphasize that even though we initialize Wally with MulPIR, it is compatible with any recent efficient PIR schemes
Each server bucket consisting of $N/B$ entries represented as $\sqrt{N/B}\times\sqrt{N/B}$ matrix $\mathbf{D}$. MulPIR uses a polynomial version of SHE, where each plaintext is a polynomial. For this explanation, we assume that each entry has the same size as SHE plaintext. If entries are larger than a plaintext, then we split each entry across multiple plaintexts, and if entries are smaller, then we pack multiple entries into each plaintext polynomial when applicable. A client query in MulPIR is a single ciphertext $\hat{q}$ encrypting row and column indices. Given a query, the server does the following.
@@
Section B discusses various optimizations we implemented for MulPIR. Overall, our optimizations resulted in a three-fold reduction in server computation and a three-fold reduction in request size.
@@
Output: Processed semantic $D_{\text{sem}}$,
Clusters centroids $c$.
@@ 5 Wally Protocol
In this section, we describe the complete Wally protocol for a single epoch in semantic search. We assume that $U$ honest clients are present to query within that epoch. Due to the space limitation, we only describe the semantic search protocol in detail, but keyword search follows a similar protocol with changes mentioned in Section B.
@@
The server uses Algorithm 1 to encode the database with $N$ embeddings into $K*d$ vectorized plaintexts. Specifically, the algorithm divides the input embeddings into $K$ clusters using K-means clustering. The algorithm then iterates over each cluster separately. Each cluster consists of $|C[i]|/d$ column-wise $d*d$ slices. The algorithm then fills $d$ plaintext vectors for each cluster. In a plaintext vector $j$, the algorithm packs $j$’th diagonal of all $|C[i]|/d$ matrices. Once $d$ plaintext vectors are generated, the algorithm iterates over a group of $g=\sqrt{d}$ plaintexts, rotating each group in increasing multiples of $-g$. These rotations are required for the baby-step giant step (BSGS) optimization [27]. Note that these rotations are performed on plaintexts, so their cost are negligible.
@@
The algorithm, then permutes the real and fake queries list $\mathbf{Q}$ and generates a random schedule for the queries. To generate the schedule $\mathbf{S}$, for each query $i\in Q$, the algorithm picks a random slot $t$ independently and appends $i$ to $S[t]$. In other words, the schedule $\mathbf{S}$ maps slot indices to the query indices. At a particular time slot $t$ within an epoch, the client will make queries (independently) indicated by $S[t]$.
@@
The server uses Algorithm 4 to generate a response for every query received. The algorithm first aligns the query ciphertext $\mathsf{ct}$. That is, first it copies the query ciphertext $g$ times and then rotates each copy left in increments of one, in total $g$ rotations. Then, the algorithm performs a dot-product between rotated query ciphertexts and the plaintext vectors for cluster $\mathsf{id}$. Recall that we ensure each cluster is encoded into $d$ plaintexts. The algorithm iterates in groups of $g$ plaintexts; within each group, multiply $i$-th plaintext with the $i$-th copy of the rotated query and sum the resulting ciphertexts. These iterations yield $h=\sqrt{d}$ ciphertexts. After that, the algorithm iterates over the $h$ resulting ciphertexts, rotating right each in increments of $g$ and summing all of them into a single ciphertext $R$. This step requires a total of $h$ rotations. The algorithm returns $R$ and cluster $\mathsf{id}$’s metadata.
@@ 6.1 DP Security in the Central Model
We define the curator as Algorithm $M$ that takes input from a database $X$ of requests from $N$ clients and outputs a noisy histogram. We prove the theorem by investigating the affected buckets when we change the input $X$ by replacing a client with another.
@@ 6.2 Generating view of the server
We use infinite divisibility of $\mathcal{NB}$ distribution to produce a standard hybrid argument in cryptography.
@@
Hybrid 1: In this hybrid, we replace all the clients with a single simulator. The input of a simulator is a list $R=\{R_{u}\}_{u}$, where $R_{u}$ is real queries for $u$-th client. The simulator then simulates each client $u$ with input $R_{u}$ as in Hybrid 0.
As the simulator internally runs each client $u$ as in Hybrid 0 with input $R_{u}$. From the adversary’s point of view, the distribution of requests is the same as Hybrid 0.
@@
    In Hybrid 1, each simulated client generates a schedule for its queries independently. While in Hybrid 2, the simulator generates a schedule for queries.
@@
Note that negative binomial distribution is infinite divisible because $\mathcal{NB}(r,p)\ \sum_{i=1}^{U}\mathcal{NB}(r/U,p)$. Thus, the distribution of per cluster fake queries per cluster in Hybrid 2 (Line 2) is the same as in Hybrid 1. Observe that a random slot is picked for each query independently in both Hybrids. Therefore, the query schedule generated in both Hybrids is indistinguishable.
@@ 7.1 Implementation
We have implemented Wally in Swift atop the Apple swift-homomorphic-encryption library [7]. Following the library, we use BFV [21] to implement the SHE parts Wally. However, our protocol could easily be implemented using other SHE schemes such as BGV [11].
@@
The choice of BFV parameters provides a trade-off between computation and communication in Wally. At a high level, larger parameters provide better computation but result in high request size. We tried multiple sets of parameters and picked the one that provided the best performance. We set polynomial modulus $n=4096$, ciphertext modulus $\log_{2}{Q}=83$, and plaintext modulus $\log_{2}{t}=15$ bits. These parameters allow a precision of six bits. To increase the precision, we use plaintext CRT optimization described in Section B. Recall that this optimization’s key benefit is increasing precision without increasing the evaluation key size (which is a dominating factor in each query). For this we set first plaintext modulus $\log_{2}{t_{0}}=15$ bits and second plaintext $\log_{2}{t_{1}}=16$. This gives us a precision of 13 bits.
Differetnial privacy parameters
@@ 7.2 Performance Evaluation
queries 39 mins 44 mins 2.8 years
@@
Table 2 describes our main experiments compared to Tiptoe. The server database consists of 35 million entries divided into around 8500 clusters. We keep each cluster size less than the number of plaintext slots in each BFV ciphertext, $n=4096$. We assume each client makes $\Delta=2$ queries per epoch for these experiments. Moreover, there are four million honest clients present in an epoch. This results in eight million real queries in an epoch.
@@
where $K$ is the total number of clusters and $U$ is the total number of clients. For our selected parameters, for each real query, the client is required to make just under two fake queries. Therefore, the total number of queries within an epoch is 24 million. In Table 2, we included fake query overhead in query request, response and total computation in Wally . Epoch size in Wally equals the total time it takes to process all the queries (real and fake).
@@
In other words, Wally trades off scalability for privacy.
Wally has significantly better performance than Tiptoe across all the metrics. Specifically, the query size (including overhead due to fake queries) in Wally 19-23 times smaller than in Tiptoe. This is because the query in Wally only covers a single cluster, while Tiptoe covers all clusters. Similarly, the response size is 3.1-3.6 smaller. Wallyś efficient server computation makes it an ideal candidate for practical deployments in comparison to Tiptoe. For example, to process eight million queries, the server with 320 cores requires around 39 minutes in Wally . In Tiptoe, the same server would require more than two years to process these queries, which is unsuitable for most practical search services.
@@
Table 3 shows how each query grows as we increase precision, i.e., as we increase the BFV parameters using our plaintext CRT optimization described in Section B. The six bit precision parameters are the same as we used in Table 2 while the $13$ bit parameters use the plaintext CRT on $\log_{2}(t_{1})=15.3$ and $\log_{2}(t_{2})=16$, $\log_{2}(t)=\log_{2}(t_{1}t_{2})=31.3$. Lastly, the 29 precision parameters use two $n=8192$ BFV parameter sets with $\log_{2}t_{1}=29$ and $\log_{2}t_{2}=30$.
@@ 7.3 Keyword Query Performance
Table 5 shows single-threaded keyword PIR server performance, measured on an Intel Xeon w3-2423 CPU with 32GB of RAM using swift-homomorphic-encryption [7]. Each entry used a BFV parameter set of $n=4096$, $\log_{2}Q=83$ with three ciphertext moduli, and $\log_{2}t=6$ using MulPIR [4]. Each response size was 27KB and the query sizes were 368 KB, 539 KB, 539 KB, and 624 KB for each respective row in the table. Note, however, that each dataset size requires manual tuning for optimal encryption and cuckoo hashing parameters.
Database entries Entry Size (Bytes) Time (ms)
100k 10 166
100k 100 465
1 million 10 565
1 million 100 1700
Table 5: Keyword-PIR, or exact query, server run times for Wally. These times include the optimizations in Section B.
@@ 8 Conclusion
9 Open science and ethics discussion
9.1 Open Science Policy
We have ensured open science policies by taking the following steps: We develop our system using an open-source library. The paper discusses and provides all the research techniques we used to develop the end system. We provide open access to the paper by sharing it on Arxiv Once accepted, we will reveal part of the code that will allow reproducibility. The entire protocol is a simple extension of the code we will open source, but we cannot provide a full implementation of the end-to-end deployed system due to our company’s IP policy.
9.2 Ethics
This paper complies with all the ethics guidelines. Specifically, we do not use private data or code without developers’ permission or run experiments with people. Our protocol protects user privacy in search systems without affecting the system’s usability; therefore, it benefits all parties. We also took essential disclosures from our employer to share the results given in our paper.
Appendix A Related Work
Protocols like Tor [16] and iCloud Private Relay [6] do not provide query privacy. Further, the server can learn the distribution of queries to each cluster when these protocols are applied to clustered databases.
Recently, there has been a number of works making private search engines using homomorphic encryption, and potentially PIR, in order to obtain full cryptographic obliviousness. Tiptoe [29], the most recent and performant, has been discussed above in-detail. Other, less efficient solutions, include Coeus [1] as well as Hers [20], with the latter providing privacy for the server’s database as well. Adjusting Wally to provide security for the server would decrease performance since we would need to add circuit privacy to the server’s SHE computation, raising the RLWE parameters to the next largest ring dimension $n$ for noise flooding[8, 35], and including a client public key in each query [23].
Appendix B Optimizations
Here we present our optimizations for both use cases: approximate nearest neighbor search and exact retrieval (PIR). Our optimizations revolve around the [10, 21] and BGV [11] SHE schemes. Further, we describe several optimizations to the MulPIR scheme [4]. First, we describe the standard RNS optimization in SHE [24] and a method to compress ciphertexts by dropping LSBs as described in Cheetah [31].
@@ Appendix B Optimizations
We have correctness as long as $\|e+e_{drop}\|_{\infty}<Q^{\prime}/2t$. Concretely, we are able to drop around $8$kB from the response size in approximate nearest neighbors search to a response size of $47$kB when combined our BFV plaintext CRT optimization described in the next subsection.
B.1 Approximate Nearest Neighbors Optimizations
@@
In general, the plaintext CRT is preferable to increasing the ring dimension since the evaluation key sizes dominate the ciphertext size: six polynomials versus 2 polynomials. This is true in the bandwidth-optimal key-switching strategy for hybrid GHS key-switching. Another advantage of the plaintext CRT is that it allows for very high precision computations which do not wrap around mod $t$. One limitation of this optimization is that we run out of NTT-friendly plaintext primes for small parameter sets, like $n=4096$ or $2048$.
B.2 PIR Optimizations
@@
In general, the cost of a ciphertext-ciphertext multiplication is much higher than a ciphertext rotation and both are much higher than addition or plaintext multiplication. For example, SEAL [13] on an AWS t2.2xlarge AWS instance has ciphertext addition and plaintext multiplication as less than $100\mu s$ while ciphertext rotation is $3.6ms$ and ciphertext multiplication is $12.1ms$ [37].
@@
We apply the lazy rescaling technique of Kim et al [32] in MulPIR’s last step, the inner product between ciphertexts $\mathbf{r}^{T}\mathbf{a}$. The main idea here is that BFV multiplications first multiply polynomials over the integers, then scale and round back to integers modulo $Q$. Kim et al [32, Appendix F] noticed that one can multiply over the integers, add over the integers, then scale and round after the additions, from the dimension number of scaling operations to one. Lazy rescaling yielded a $15-20\%$ improvement in our MulPIR implementation.
@@
This approach results in $d_{1}\cdot d_{2}\cdot n$ ciphertext-plaintext multiplications and $d_{2}\cdot t$ ciphertext-ciphertext multiplications, where $t$ is the number of plaintexts used to encode one entry. The number of ciphertext-ciphertext multiplication grows linearly with $t$ and remains a concretely small value when entries are not large. But it will blow up with extremely large entries (e.g., when they are high-resolution photos).
@@
The outer-product takes $d_{1}\cdot d_{2}$ ciphertext-ciphertext mutliplications to compute, which is independent of the entry size. While the inner products still takes $d_{1}\cdot d_{2}\cdot n$ ciphertext-plaintext multiplications. Therefore, this optimization can reduce the computation overhead when the entries size $t$ is larger than the first dimension size $d_{1}$.
@@
We add a conditional, lazy modular reduce to Harvey’s NTT [28]. The main idea is that we only reduce during the butterfly if the integers would exceeds the machine word size. This saw a $20-25\%$ improvement in forward NTT times and $15-20\%$ improvement in inverse NTT times for $61$-bit primes. Note, we saw no improvement on primes near $64$-bits.

### v4

@@ 4.2 SHE details in keyword search
At a high level, for each bucket independently, the server uses the cuckoo hashing to map the $N^{\prime}$ entries into a table of size $O(N^{\prime})$ using two random hash functions. The guarantee is that if a keyword is present in the database, it must be at bucket index $h_{1}(\textit{keyword})$ or $h_{2}(\textit{keyword})$ in the table. The client then uses index PIR to privately fetch entries at these indices and locally find the index containing the entry and associated data. Wally uses MulPIR [4] as the underlying index PIR, an efficient single-server PIR scheme based on SHE. We emphasize that even though we initialize Wally with MulPIR, it is compatible with any recent efficient PIR scheme.
Each server bucket consisting of $N/B$ entries is represented as a $\sqrt{N/B}\times\sqrt{N/B}$ matrix $\mathbf{D}$. MulPIR uses a polynomial version of SHE, where each plaintext is a polynomial. For this explanation, we assume that each entry has the same size as SHE plaintext. If entries are larger than a plaintext, then we split each entry across multiple plaintexts, and if entries are smaller, then we pack multiple entries into each plaintext polynomial when applicable. A client query in MulPIR is a single ciphertext $\hat{q}$ encrypting row and column indices. Given a query, the server does the following.
@@
Section A discusses various optimizations we implemented for MulPIR. Overall, our optimizations resulted in a three-fold reduction in server computation and a three-fold reduction in request size.
@@
Output: Processed semantic $D_{\text{sem}}$, cluster centroids $c$.
@@ 5 Wally Protocol
In this section, we describe the complete Wally protocol for a single epoch in semantic search. We assume that $U$ honest clients are present to query within that epoch. Due to the space limitation, we only describe the semantic search protocol in detail, but keyword search follows a similar protocol with changes mentioned in Section A.
@@
The server uses Algorithm 1 to encode the database with $N$ embeddings into $K*d$ vectorized plaintexts. Specifically, the algorithm divides the input embeddings into $K$ clusters using K-means clustering. The algorithm then iterates over each cluster separately. Each cluster consists of $|C[i]|/d$ column-wise $d\times d$ slices. The algorithm then fills $d$ plaintext vectors for each cluster. In a plaintext vector $j$, the algorithm packs $j$’th diagonal of all $|C[i]|/d$ matrices. Once $d$ plaintext vectors are generated, the algorithm iterates over a group of $g=\sqrt{d}$ plaintexts, rotating each group in increasing multiples of $-g$. These rotations are required for the baby-step giant step (BSGS) optimization [24]. Note that these rotations are performed on plaintexts, so their cost are negligible.
@@
The algorithm then permutes the real and fake queries list $\mathbf{Q}$ and generates a random schedule for the queries. To generate the schedule $\mathbf{S}$, for each query $i\in Q$, the algorithm picks a random slot $t$ independently and appends $i$ to $S[t]$. In other words, the schedule $\mathbf{S}$ maps slot indices to the query indices. At a particular time slot $t$ within an epoch, the client will make queries (independently) indicated by $S[t]$.
@@
The server uses Algorithm 4 to generate a response for every query received. The algorithm first aligns the query ciphertext $\mathsf{ct}$. That is, first it copies the query ciphertext $g$ times and then rotates each copy left in increments of one, in total $g$ rotations. Then, the algorithm performs a dot-product between rotated query ciphertexts and the plaintext vectors for cluster $\mathsf{id}$. Recall that we ensure each cluster is encoded into $d$ plaintexts. The algorithm iterates in groups of $g$ plaintexts; within each group, multiply the $i$-th plaintext with the $i$-th copy of the rotated query and sum the resulting ciphertexts. These iterations yield $h=\sqrt{d}$ ciphertexts. After that, the algorithm iterates over the $h$ resulting ciphertexts, rotating right each in increments of $g$ and summing all of them into a single ciphertext $R$. This step requires a total of $h$ rotations. The algorithm returns $R$ and cluster $\mathsf{id}$’s metadata.
@@ 6.1 DP Security in the Central Model
We define the curator as Algorithm $M$ that takes input from a database $X$ of requests from $N$ clients and outputs a noisy histogram. We prove the theorem by investigating the affected buckets when we change the input $X$ by replacing one client with another.
@@ 6.2 Generating view of the server
We use infinite divisibility of $\mathcal{NB}$ distribution to produce a standard cryptography hybrid argument.
@@
Hybrid 1: In this hybrid, we replace all the clients with a single simulator. The input of a simulator is a list $R=\{R_{u}\}_{u}$, where $R_{u}$ are the real queries for $u$-th client. The simulator then simulates each client $u$ with input $R_{u}$ as in Hybrid 0.
As the simulator internally runs each client $u$ as in Hybrid 0 with input $R_{u}$, from the adversary’s point of view, the distribution of requests is the same as Hybrid 0.
@@
    In Hybrid 1, each simulated client generates a schedule for its queries independently, while in Hybrid 2, the simulator generates a schedule for queries.
@@
Note that negative binomial distribution is infinite divisible because $\mathcal{NB}(r,p)=\sum_{i=1}^{U}\mathcal{NB}(r/U,p)$. Thus, the distribution of per cluster fake queries per cluster in Hybrid 2 (Line 2) is the same as in Hybrid 1. Observe that a random slot is picked for each query independently in both Hybrids. Therefore, the query schedule generated in both Hybrids is indistinguishable.
@@ 7.1 Implementation
We have implemented Wally in Swift atop the Apple swift-homomorphic-encryption library [6]. Following the library, we use BFV [18] to implement the SHE parts of Wally. However, our protocol could easily be implemented using other SHE schemes such as BGV [10].
@@
The choice of BFV parameters provides a trade-off between computation and communication in Wally. At a high level, larger parameters provide better computation but result in high request size. We tried multiple sets of parameters and picked the one that provided the best performance. We set polynomial modulus $n=4096$, ciphertext modulus $\log_{2}{Q}=83$, and plaintext modulus $\log_{2}{t}=15$ bits. These parameters allow a precision of six bits. To increase the precision, we use plaintext CRT optimization described in Section A. This optimization’s key benefit is increasing precision without increasing the evaluation key size (which is a dominating factor in each query). For this we set first plaintext modulus $\log_{2}{t_{0}}=15$ bits and second plaintext $\log_{2}{t_{1}}=16$. This gives us a precision of 13 bits.
Differential privacy parameters
@@ 7.2 Performance Evaluation
queries 117 mins 139 mins 2.8 years
@@
Table 2 describes our main experiments compared to Tiptoe. For an end-to-end system, we estimate the network time on a 30Mbps channel. This comes out to 74ms for transmitting a query and response over the network for the small precision setting (first column in Tables 2 and 3).
The server database consists of 35 million entries divided into around 8500 clusters. We keep each cluster size less than the number of plaintext slots in each BFV ciphertext, $n=4096$. We assume each client makes $\Delta=2$ queries per epoch for these experiments. Moreover, there are four million honest clients present in an epoch. This results in eight million real queries in an epoch.
@@
where $K$ is the total number of clusters and $U$ is the total number of clients. For our selected parameters, for each real query, the client is required to make just under two fake queries. Therefore, the total number of queries within an epoch is 24 million. In Table 2, we included fake query overhead in query request, response and total computation plus estimated network time in Wally. The epoch size in Wally is the total time it takes to process all the queries (real and fake).
@@
In other words, Wally trades off privacy guarantees for scalability.
Wally has significantly better performance than Tiptoe across all the metrics. Specifically, the query size (including overhead due to fake queries) in Wally 19-23 times smaller than in Tiptoe. This is because the query in Wally only covers a single cluster, while Tiptoe covers all clusters. Similarly, the response size is 3.1-3.6 smaller. Wally’s efficient server computation makes it an ideal candidate for practical deployments in comparison to Tiptoe. For example, to process eight million queries, the server with 320 cores requires around 117 minutes in Wally. In Tiptoe, the same server would require more than two years to process these queries, which is unsuitable for most practical search services.
@@
Table 3 shows how each query grows as we increase precision, i.e., as we increase the BFV parameters using our plaintext CRT optimization described in Section A. The six bit precision parameters are the same as we used in Table 2 while the $13$ bit parameters use the plaintext CRT on $\log_{2}(t_{1})=15.3$ and $\log_{2}(t_{2})=16$, $\log_{2}(t)=\log_{2}(t_{1}t_{2})=31.3$. Lastly, the 29 precision parameters use two $n=8192$ BFV parameter sets with $\log_{2}t_{1}=29$ and $\log_{2}t_{2}=30$.
@@ 7.3 Keyword Query Performance
Table 5 shows single-threaded keyword PIR server performance, measured on an Intel Xeon w3-2423 CPU with 32GB of RAM using swift-homomorphic-encryption [6]. Each entry used a BFV parameter set of $n=4096$, $\log_{2}Q=83$ with three ciphertext moduli, and $\log_{2}t=6$ using MulPIR [4]. Each response size was 27KB. Note, however, that each dataset size requires manual tuning for optimal encryption and cuckoo hashing parameters.
Database entries Entry Size (Bytes) Query size (KB) Time (ms)
100k 10 368 166
100k 100 539 465
1 million 10 539 565
1 million 100 624 1700
Table 5: Keyword-PIR, or exact query, server run times for Wally. These times include the optimizations in Section A.
8 Related Work
Anonymization networks alone do not provide query privacy. Further, the server can learn the distribution of queries to each cluster when these protocols are applied to clustered databases.
Recently, there has been a number of works making private search engines using homomorphic encryption, and potentially PIR, in order to obtain full cryptographic obliviousness. Tiptoe [26], the most recent and performant, has been discussed above in-detail. Other less efficient solutions include Coeus [1] as well as Hers [17], with the latter providing privacy for the server’s database as well. Adjusting Wally to provide security for the server would decrease performance since we would need to add circuit privacy to the server’s SHE computation, raising the RLWE parameters to the next largest ring dimension $n$ for noise flooding[7, 32], and including a client public key in each query [20].
@@ 9 Conclusion
Appendix A Optimizations
Here we present our optimizations for both use cases: approximate nearest neighbor search and exact retrieval (PIR). Our optimizations revolve around the BFV [9, 18] and BGV [10] SHE schemes. Further, we describe several optimizations to the MulPIR scheme [4]. First, we describe the standard RNS optimization in SHE [21] and a method to compress ciphertexts by dropping LSBs as described in Cheetah [28].
@@ Appendix A Optimizations
We have correctness as long as $\|e+e_{drop}\|_{\infty}<Q^{\prime}/2t$. Concretely, we are able to drop around $8$kB from the response size in approximate nearest neighbors search to a response size of $47$kB when combined with our BFV plaintext CRT optimization described in the next subsection.
A.1 Approximate Nearest Neighbors Optimizations
@@
In general, plaintext CRT is preferable to increasing the ring dimension since the evaluation key sizes dominate the ciphertext size: 6 polynomials versus 2 polynomials. This is true in the bandwidth-optimal key-switching strategy for hybrid GHS key-switching. Another advantage of the plaintext CRT is that it allows for very high precision computations which do not wrap around mod $t$. One limitation of this optimization is that we run out of NTT-friendly plaintext primes for small parameter sets, like $n=4096$ or $2048$.
A.2 PIR Optimizations
@@
In general, the cost of a ciphertext-ciphertext multiplication is much higher than a ciphertext rotation and both are much higher than addition or plaintext multiplication. For example, Table 1 shows ciphertext addition and plaintext multiplication as less than $100\mu s$ while ciphertext rotation is $0.5ms$ and ciphertext multiplication is $2.5ms$.
@@
We apply the lazy rescaling technique of Kim et al [29] in MulPIR’s last step, the inner product between ciphertexts $\mathbf{r}^{T}\mathbf{a}$. The main idea here is that BFV multiplications first multiply polynomials over the integers, then scale and round back to integers modulo $Q$. Kim et al [29, Appendix F] noticed that one can multiply over the integers, add over the integers, then scale and round after the additions, from the dimension number of scaling operations to one. Lazy rescaling yielded a $15-20\%$ improvement in our MulPIR implementation.
@@
This approach results in $d_{1}\cdot d_{2}\cdot n$ ciphertext-plaintext multiplications and $d_{2}\cdot t$ ciphertext-ciphertext multiplications, where $t$ is the number of plaintexts used to encode one entry. The number of ciphertext-ciphertext multiplications grows linearly with $t$ and remains a concretely small value when entries are not large. However, it will blow up with extremely large entries (e.g., when they are high-resolution photos).
@@
The outer-product takes $d_{1}\cdot d_{2}$ ciphertext-ciphertext multiplications to compute, which is independent of the entry size, while the inner products still take $d_{1}\cdot d_{2}\cdot n$ ciphertext-plaintext multiplications. Therefore, this optimization can reduce the computation overhead when the entry size ($t$ times the plaintext size) is larger than the first dimension size $d_{1}$.
@@
We add a conditional, lazy modular reduction to Harvey’s NTT [25]. The main idea is that we only reduce during the butterfly if the integers would exceed the machine word size. This saw a $20-25\%$ improvement in forward NTT times and $15-20\%$ improvement in inverse NTT times for $61$-bit primes. Note, we saw no improvement on primes near $64$-bits.